# WeFT: Weighted Entropy-driven Fine-Tuning for dLLMs

### v1

@@ A.2 Proof of the Main Theorem
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})(s_{\theta}(x,t)_{x_{t}^{\prime}}-\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}+C)dt$
@@
$s_{\theta}(x,t)_{x_{t}^{\prime}}\approx\frac{p_{t}(x_{t}^{1}\dots x^{\prime}{}^{i}_{t}\dots x_{t}^{d})}{p_{t}(x_{t}^{1}\dots x^{i}_{t}\dots x_{t}^{d})}$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})s_{\theta}(x,t)_{x_{t}^{\prime}}dt=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}\beta_{i}f(t)\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}dt$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})(-\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}})dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M],j\neq[M]}-f(t)\beta_{i}\mathbf{1}[x_{0}^{i}=j]\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}dt$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})K(\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})})$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M],j\neq[M]}\beta_{i}f(t)K(\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}})\bar{f}(t))}\mathbf{1}[x_{0}^{i}=j])dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}\beta_{i}f(t)\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}})\bar{f}(t))}(\log\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}})\bar{f}(t))}-1)dt$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}-\frac{\beta_{i}f(t)\exp(-\beta_{i}\bar{f}(t))}{1-\exp(-\beta_{i}\bar{f}(t))}\log(\frac{\exp(-\beta_{i}\bar{f}(t))}{1-\exp(-\beta_{i}\bar{f}(t))}q_{\theta}(x_{0}^{i}|x_{t}^{\text{unmasked}}))dt$
@@
$t_{i}=1-\exp\!\big(-\beta_{x_{t}^{i}}\,\bar{f}(t)\big),dt_{i}=\beta_{x_{t}^{i}}f(t)\exp\!\big(-\beta_{x_{t}^{i}}\,\bar{f}(t)\big)$
the loss can be rewritten as:
$\sum_{i}\int_{0}^{1}\mathbb{E}_{x\sim p_{t|0}(x_{t}|x_{0})}[-\frac{1}{t_{i}}\mathbf{1}[x_{t}^{i}=[\mathbf{M}]]\log(q_{\theta}(x_{0}^{i}|x_{t}^{\text{unmasked}})]dt_{i}+C$
@@
$t_{i}=1-(1-t)^{\frac{\beta_{x^{i}}}{\beta_{\text{ref}}}}.$ (21)
@@
Appendix B Hyperparameters
B.1 Training Hyperparameters
Our experiments are conducted with a learning rate of $1\times 10^{-5}$ and a total of 20 training epochs. We apply gradient accumulation with $4$ steps and used an effective per-device training batch size of $1$. To reduce memory cost, we employ LoRA with rank $r=128$, scaling factor $\alpha=256$, and a dropout rate of $0.05$. All hyperparameters, including model architecture, optimization, LoRA configuration, and training settings, are listed in Table 6.
Category Hyperparameter = Value
Model Architecture d_model = 4096; n_layers = 32; n_heads = 32; n_kv_heads = 32;
vocab_size=126464\ =126464; rope == True; rope_theta=500000\ =500000;
pad_token_id=126081\ =126081; mask_token_id=126336\ =126336;
[…]
Optimization optimizer=\ =\ adamw_torch; weight_decay=0.1\ =0.1;
learning_rate = 1e-5; lr_scheduler_type = linear;
adam_beta1 = 0.9; adam_beta2 = 0.999; adam_epsilon = 1e-8;
init_fn=\ =\ mitchell; init_std=0.02\ =0.02; max_grad_norm=1\ =1
target_modules == [”q_proj”, ”k_proj”, ”v_proj”]
gradient_accumulation_steps=4\ =4; precision = bf16;
max_sequence_length = 4096; seed =42\ =42
Table 6: Hyperparameters used in training.
B.2 Evaluation Hyperparameters
In the evaluation, we set the block length to 32. For the two smaller datasets, Sudoku and Countdown, the generation length is set to 512; for the two larger datasets, MATH500 and GSM8K, the generation length is set to 256. The number of diffusion steps is always set to half of the generation length, and we use the low-confidence remasking method.
Appendix C Algorithm Pipeline
In Algorithm 1, we provide a detailed description of the implementation of WeFT.
Output: Loss $\mathcal{L}$
Function EstimateRate($\mathcal{M},\mathcal{I},l,M,m$):
@@ Appendix C Algorithm Pipeline
 $\beta_{i}\leftarrow(1-M^{i})\sqrt{H_{i}}$;
@@
$\mathcal{L}\leftarrow\frac{\sum_{i}\mathbf{1}_{\{M^{\prime i}=0\}}\frac{1}{t_{i}}\text{CrossEntropy}(z_{i},l_{i})}{\sum_{i}\mathbf{1}_{\{M^{\prime i}=0\}}}$;
@@
Algorithm 1 WeFT: Weighted Entropy-driven Fine-Tuning

### v2

@@ A.2 Proof of the Main Theorem
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}$
$\quad\times\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})\Big(s_{\theta}(x,t)_{x_{t}^{\prime}}$
$\quad-\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}+C\Big)dt$
@@
$s_{\theta}(x,t)_{x_{t}^{\prime}}$ $\approx p_{t}(x_{t}^{1}\dots x^{\prime}{}^{i}_{t}\dots x_{t}^{d})$
$\quad\times\frac{1}{p_{t}(x_{t}^{1}\dots x^{i}_{t}\dots x_{t}^{d})}$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})s_{\theta}(x,t)_{x_{t}^{\prime}}dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}\beta_{i}f(t)$
$\quad\times\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}dt$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})$
$\quad\times\left(-\frac{p_{t|0}(x_{t}^{\prime}|x_{0})}{p_{t|0}(x_{t}|x_{0})}\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}\right)dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M],j\neq[M]}-f(t)\beta_{i}$
$\quad\times\mathbf{1}[x_{0}^{i}=j]\log s_{\theta}(x_{t},t)_{x_{t}^{\prime}}dt$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}\mid x_{0})}\sum_{x_{t}^{\prime}\neq x_{t}}Q_{t}(x_{t}^{\prime},x_{t})$
$\quad\times K\!\left(\frac{p_{t|0}(x_{t}^{\prime}\mid x_{0})}{p_{t|0}(x_{t}\mid x_{0})}\right)\,dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}\mid x_{0})}\sum_{x_{t}^{i}=[M],j\neq[M]}\beta_{i}f(t)$
$\quad\times K\!\left(\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}})\bar{f}(t)}\mathbf{1}[x_{0}^{i}=j]\right)dt$
$=\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}\mid x_{0})}\sum_{x_{t}^{i}=[M]}\beta_{i}f(t)$
$\quad\times\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}$
$\quad\times\left(\log\frac{\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}{1-\exp(-\beta_{x_{t}^{i}}\bar{f}(t))}-1\right)dt$
@@
$\int_{0}^{T}\mathbb{E}_{x_{t}\sim p_{t|0}(x_{t}|x_{0})}\sum_{x_{t}^{i}=[M]}-\frac{\beta_{i}f(t)\exp(-\beta_{i}\bar{f}(t))}{1-\exp(-\beta_{i}\bar{f}(t))}$
$\quad\times\log\left(\frac{\exp(-\beta_{i}\bar{f}(t))}{1-\exp(-\beta_{i}\bar{f}(t))}q_{\theta}(x_{0}^{i}|x_{t}^{\text{unmasked}})\right)dt$
@@
$t_{i}=1-\exp\!\big(-\beta_{x_{t}^{i}}\,\bar{f}(t)\big),$
we obtain
$dt_{i}=\beta_{x_{t}^{i}}f(t)\exp\!\big(-\beta_{x_{t}^{i}}\,\bar{f}(t)\big)dt$
and the loss can be rewritten as:
$\sum_{i}\int_{0}^{1}\mathbb{E}_{x\sim p_{t|0}(x_{t}|x_{0})}\Big[-\frac{1}{t_{i}}\mathbf{1}[x_{t}^{i}=[M]]$
$\quad\times\log(q_{\theta}(x_{0}^{i}|x_{t}^{\text{unmasked}}))\Big]dt_{i}+C$
@@
$t_{i}$ $=1-\exp(-\beta_{x^{i}}\,\bar{f}(t))$
$=1-\exp(-\beta_{x^{i}}(\frac{1}{\beta_{ref}}\log(1-t)))$
$=1-(1-t)^{\frac{\beta_{x^{i}}}{\beta_{\text{ref}}}}$
@@
Appendix B Algorithm Pipeline
In Algorithm 1 and Algorithm 2, we provide a detailed description of the implementation of GIFT.
Input: model $\mathcal{M}$, input_ids $\mathcal{I}$, mask_token_id $m$
Output: Masking parameter $\beta$
Function EstimateRate($\mathcal{M},\mathcal{I},M,m$):
@@ Appendix B Algorithm Pipeline
 $\beta_{i}\leftarrow M^{i}\sqrt{H_{i}}$;
@@
Algorithm 1 Creating Betas from Entropy
Output: Loss $\mathcal{L}$
@@
$\mathcal{I^{\prime}}\leftarrow(1-M^{\prime})\cdot\mathcal{I}+M^{\prime}\cdot m$;
$z\leftarrow\mathcal{M}(\mathcal{I^{\prime}}).\text{logits}$;
$S\leftarrow\sum_{i}\mathbf{1}_{\{M^{\prime i}=1\}}\frac{1}{t_{i}}\cdot$ $\text{CrossEntropy}(z_{i},l_{i})$;
$N\leftarrow\sum_{i}\mathbf{1}_{\{M^{\prime i}=1\}}$;
$\mathcal{L}\leftarrow S/N$;
@@
Algorithm 2 Weighted Entropy-driven Fine-Tuning
Category Hyperparameter = Value
Optimization optimizer=\ =\ adamw_torch; weight_decay=0.1\ =0.1;
learning_rate = 1e-5; lr_scheduler_type = linear;
adam_beta1 = 0.9; adam_beta2 = 0.999; adam_epsilon = 1e-8;
init_fn=\ =\ mitchell; init_std=0.02\ =0.02; max_grad_norm=1\ =1
target_modules == ["q_proj", "k_proj", "v_proj"]
gradient_accumulation_steps=4\ =4; precision = bf16;
max_sequence_length = 4096; seed =42\ =42
Table 6: Hyperparameters used in LoRA training.
Category Hyperparameter = Value
Optimization optimizer=\ =\ adamw_torch; weight_decay=0\ =0;
learning_rate = 2e-5; lr_scheduler_type = cosine;
adam_beta1 = 0.9; adam_beta2 = 0.999; adam_epsilon = 1e-8;
init_fn=\ =\ mitchell; init_std=0.02\ =0.02; max_grad_norm=1\ =1
Training num_train_epochs=5\ =5; per_device_train_batch_size=4\ =4;
gradient_accumulation_steps=1\ =1; precision = bf16;
max_sequence_length = 4096; seed =42\ =42
Table 7: Hyperparameters used in full parameter training.
Appendix C Hyperparameters and Detailed Settings
C.1 Training Hyperparameters and Settings
Our training is divided into two settings: LoRA fine-tuning and full parameter fine-tuning. LoRA fine-tuning is conducted on 4 H100 GPUs, while full parameter fine-tuning is conducted on 8 H100 GPUs.
For LoRA fine-tuning, experiments are conducted with a learning rate of $1\times 10^{-5}$ and a total of 20 training epochs. We apply gradient accumulation with $4$ steps and used an effective per-device training batch size of $1$. To reduce memory cost, we employ LoRA with rank $r=128$, scaling factor $\alpha=256$, and a dropout rate of $0.05$. All hyperparameters, including model architecture, optimization, LoRA configuration, and training settings, are listed in Table 6.
For full parameter fine-tuning, we train for 5 epochs with a learning rate of $2\times 10^{-5}$. Other settings are detailed in Table 7.
C.2 Evaluation Hyperparameters and Settings
In the evaluation, we set the block length to 32. For the two smaller datasets, Sudoku and Countdown, the generation length is set to 512; for the two larger datasets, MATH500 and GSM8K, the generation length is set to 256. The number of diffusion steps is always set to half of the generation length, and we use the low-confidence remasking method. Regarding the specific partitioning of validation sets for these four datasets, we follow the settings in wd1 (Tang et al., 2025).
Appendix D Motivation for Using Entropy-based Methods as the Weight Coefficient
In this section, we provide an intuitive perspective to justify the choice of entropy as the coefficient $\beta$, as this section is intended to provide motivation rather than a rigorous proof.
In Alain et al. (2016), the authors proved that importance sampling has minimum variance when the sampling proposal is proportional to the $L_{2}$-norm of the gradient.
Recall that in a model, $p$ is generated by applying softmax to the output logits. More specifically,
$p_{k}=\frac{e^{z_{k}}}{\sum_{j}e^{z_{j}}}$
Here $z$ is the logits generated by the model.
Let $c$ be the label and $p_{c}$ be the probability of this token. Then we have
$\frac{\partial\log p_{c}}{\partial z_{k}}=\left\{\begin{array}[]{ll}-p_{k}&\text{if }k\neq c\\ 1-p_{c}&\text{if }k=c\end{array}\right$.
In the following discussion, we assume that the base model generates the correct token with relatively high probability and other tokens with relatively low probability. Therefore, $L_{2}$-norm of $\frac{\partial\log p_{c}}{\partial\mathbf{z}}$ can be approximated as $1-p_{c}$. Since $\frac{\partial\log p_{c}}{\partial\mathbf{z}}$ constitutes a key component of the loss gradient, theoretical insights suggest that selecting an importance sampling distribution that is positively correlated with this term might help reduce the variance of the loss, thereby enhancing training stability and effectiveness. We now briefly analyze why entropy can serve this purpose.
The entropy function $H(\mathbf{p})$ admits the following lower bound:
$H(\mathbf{p})$ $\geq-p_{c}\log p_{c}+(1-p_{c})$ (20)
$(\log(n-1)-\log(1-p_{c}))$
where $n$ denotes the vocabulary size. Since the term $p_{c}\log p_{c}$ satisfies
$p_{c}\log p_{c}<p_{c}(1-p_{c}),$ (21)
it’s negligible compared to $(1-p_{c})\log(n-1)$. We may safely ignore it for approximation purposes.
Moreover, we empirically observe that the variation in incorrect predictions across different samples is relatively small. Under this assumption, $H(\mathbf{p})$ is positively correlated with $(1-p_{c})$.
On the other hand, using entropy $H(\mathbf{p})$ is superior to directly using $(1-p_{c})$ in that it preserves richer information about the underlying distribution. Due to the prevalence of synonyms and semantically similar tokens, the true probability distribution is unlikely to be unimodal. Instead, it is more reasonable to expect the probability mass to be concentrated on a small subset of tokens. In this respect, entropy $H(\mathbf{p})$ captures such structural information significantly better than $(1-p_{c})$. In pratice, we employ a power function based on $\sqrt{H(\mathbf{p})}$ as the importance weighting function.
Appendix E SFT Role in Subsequent RL
Figure 3: Reward curves of models cold-started with GIFT and SFT during subsequent reinforcement learning training. The curves are smoothed using a time-weighted EMA. As shown, the model initialized with GIFT achieves higher rewards.
Reinforcement learning (RL) post-training has become a crucial stage in the training of large language models  (DeepSeek-AI, 2025). Supervised fine-tuning (SFT) cold-start plays an important role in preparing the model for subsequent RL, as SFT establishes a reliable initial policy which can then be further refined by RL objectives. Therefore, it is essential to verify whether models trained with GIFT can maintain their advantages during subsequent RL training. To this end, we fine-tune models on the s1K dataset using both GIFT and SFT, and then further train them with RL on the Sudoku dataset for comparison.
As shown in Figure 3, models initialized with GIFT achieve higher rewards throughout RL training. Moreover, as reported in Table 8, the GIFT-trained model achieves significantly better evaluation results. These findings demonstrate that models cold-started with GIFT are effective in guiding subsequent RL learning.
Model Base SFT GIFT SFT+RL GIFT+RL
Sudoku 5.5 4.6 7.8 8.9 13.3
Table 8: Evaluation results of models cold-started with GIFT and SFT during subsequent RL training.